# Mobile-Agent: Autonomous Multi-Modal Mobile Device Agent with Visual Perception

### v1

@@ Abstract
Mobile device agent based on Multimodal Large Language Models (MLLM) is becoming a popular application. In this paper, we introduce Mobile-Agent, an autonomous multi-modal mobile device agent. Mobile-Agent first leverages visual perception tools to accurately identify and locate both the visual and textual elements within the app’s front-end interface. Based on the perceived vision context, it then autonomously plans and decomposes the complex operation task, and navigates the mobile Apps through operations step by step. Different from previous solutions that rely on XML files of Apps or mobile system metadata, Mobile-Agent allows for greater adaptability across diverse mobile operating environments in a vision-centric way, thereby eliminating the necessity for system-specific customizations. To assess the performance of Mobile-Agent, we introduced Mobile-Eval, a benchmark for evaluating mobile device operations. Based on Mobile-Eval, we conducted a comprehensive evaluation of Mobile-Agent. The experimental results indicate that Mobile-Agent achieved remarkable accuracy and completion rates. Even with challenging instructions, such as multi-app operations, Mobile-Agent can still complete the requirements. Code and model will be open-sourced at https://github.com/X-PLUG/MobileAgent.
@@ 3 Experiments
Alibaba.com
Help me find caps in Alibaba.com.
Help me find caps in Alibaba.com. If the "Add to cart" is avaliable in the item information page, please add the item to my cart.
@@
Amazon Music
Search singer Jay Chou in Amazon Music.
@@
Chrome
Search result for today’s Lakers game.
@@
Gmail
Send an empty email to to {address}.
@@
Google Maps
Navigate to Hangzhou West Lake.
@@
Google Play
Download WhatsApp in Play Store.
@@
Notes
Create a new note in Notes.
@@
Settings
Turn on the dark mode.
@@
TikTok
Swipe a video about pet cat in TikTok and click a "like" for this video.
@@
YouTube
Search for videos about Stephen Curry on YouTube.
@@
Multi-App
Open the calendar and look at today’s date, then go to Notes and create a new note to write "Today is [today’s data]".

### v2

@@ Abstract
Mobile device agent based on Multimodal Large Language Models (MLLM) is becoming a popular application. In this paper, we introduce Mobile-Agent, an autonomous multi-modal mobile device agent. Mobile-Agent first leverages visual perception tools to accurately identify and locate both the visual and textual elements within the app’s front-end interface. Based on the perceived vision context, it then autonomously plans and decomposes the complex operation task, and navigates the mobile Apps through operations step by step. Different from previous solutions that rely on XML files of Apps or mobile system metadata, Mobile-Agent allows for greater adaptability across diverse mobile operating environments in a vision-centric way, thereby eliminating the necessity for system-specific customizations. To assess the performance of Mobile-Agent, we introduced Mobile-Eval, a benchmark for evaluating mobile device operations. Based on Mobile-Eval, we conducted a comprehensive evaluation of Mobile-Agent. The experimental results indicate that Mobile-Agent achieved remarkable accuracy and completion rates. Even with challenging instructions, such as multi-app operations, Mobile-Agent can still complete the requirements. Code and model are open-sourced at https://github.com/X-PLUG/MobileAgent.
@@ 3 Experiments
Alibaba.com Help me find caps in Alibaba.com.
Help me find caps in Alibaba.com. If the "Add to cart" is available in the item information page, please add the item to my cart.
@@
Amazon Music Search singer Jay Chou in Amazon Music.
@@
Chrome Search result for today’s Lakers game.
@@
Gmail Send an empty email to to {address}.
@@
Google Maps Navigate to Hangzhou West Lake.
@@
Google Play Download WhatsApp in Play Store.
@@
Notes Create a new note in Notes.
@@
Settings Turn on the dark mode.
@@
TikTok Swipe a video about pet cat in TikTok and click a "like" for this video.
@@
YouTube Search for videos about Stephen Curry on YouTube.
@@
Multi-App Open the calendar and look at today’s date, then go to Notes and create a new note to write "Today is [today’s data]".